trish:

Identify the 10th term for the sequence: 12,6,3,1.5...?

Hero:

@trish what can we divide 12 by to get 6?

trish:

@Hero 0.5

Hero:

Actually you divide 12 by 2 to get 6.

Hero:

You can multiply by 0.5 to get 6 as well.

Hero:

From my perspective, dividing by 2 is easier than multiplying by decimals.

trish:

@Hero when I do that both ways I'm getting a long decimal answer and I'm pretty sure it's not the answer

Hero:

But I understand why you did it. Neverthless it should be obvious how to get to the tenth term.

Hero:

Okay, so let's do it one step at a time.

Hero:

I'll show you how it is easy. You just have to be comfortable converting between fraction and decimal.

Hero:

So let's start with 1.5 Remember, to read 1.5 you say one AND five tenths which we can re-write in fraction form as \(1 + \dfrac{5}{10}\) but reduces to \(1 + \dfrac{1}{2}\)

Hero:

Now suppose you take \(1 + \dfrac{1}{2}\) and multiply that by \(\dfrac{1}{2}\). If you do that you get \(\dfrac{1}{2} + \dfrac{1}{4}\) which adds to \(\dfrac{3}{4}\) or \(0.75\) in decimal.

Hero:

Now that's our fifth term. Since we have \(\dfrac{3}{4}\) and need to get the sixth term, all we have to do is multiply \(\dfrac{3}{4}\) by \(\dfrac{1}{2}\) again to get the next term

trish:

@Hero as a fraction I got 3/128

Hero:

Taking another angle at this, \(1.5\) is \(\dfrac{3}{2}\) in fraction form. You can multiply that by \(\dfrac{1}{2}\) to easily get \(\dfrac{3}{4}\)

Hero:

Since \(\dfrac{3}{2} \times \dfrac{1}{2} = \dfrac{3}{4}\)

trish:

@Hero yes I understand the process but I'm hesitant that my answer is incorrect

Hero:

Okay, so the pattern is \(a_{n-1} \times \dfrac{1}{2} = a_n\) Let: \(a_1 = 12\) \(a_2 = 6\) \(a_3 = 3\) \(a_4 = \dfrac{3}{2}\)

Hero:

Then \(\begin{align*}a_5 &= a_4 \times \dfrac{1}{2} \\&= \dfrac{3}{2} \times \dfrac{1}{2} \\&= \dfrac{3}{4} \end{align*}\)

Hero:

So \(a_5 = \dfrac{3}{4}\)

Hero:

To get \(a_6\) do the same thing.

Hero:

\(a_6 = a_5 \times \dfrac{1}{2}\)

Hero:

\(a_6 = \dfrac{3}{4} \times \dfrac{1}{2}\)

trish:

@Hero thanks so much I understand it now

Hero:

Yeah but you never stated the tenth term.